# thanks: corresponding author

### v1

@@ 3 Applications: SVM Regression in Physical Systems
Building on the equivalence established between SVM kernels and Green’s functions, we now explore various applications of SVM regression in modeling physical systems, demonstrating how aligning kernel functions with corresponding Green’s functions enhances predictive accuracy. We examine cases including electrical conductivity, scattering amplitudes, anharmonic oscillator energy levels, and photonic crystals.
3.1 Computing Conductivity Using SVM Regression
Data Acquisition
For this example, we used electronic band structures, Fermi levels, and $\mathbf{k}$-points for copper, which were retrieved from the Materials Project database [29]. Each $\mathbf{k}_{i}$ corresponds to an electronic energy level $\epsilon_{i}$ in copper’s band structure.
Feature Construction
The features included group velocities for electronic states, Fermi occupation, and local conductivity contribution, with details as follows:
  1. 1.
    Velocities $v_{i}$ were computed by numerically differentiating the energy bands w.r.t. $\mathbf{k}$-points, giving the group velocity for each electronic state.
  2. 2.
    Fermi occupation $n_{F}(\epsilon_{i})$ was calculated under an assumed temperature $T$ and Fermi level $\mu$,
[…]
  3. 3.
    Local conductivity contribution $\sigma_{i}$ was derived from
    $\sigma_{i}\;=\;e^{2}\,v_{i}^{2}\,n_{F}(\epsilon_{i})\,\tau,$ (23)
    where $e$ is the electron charge, $\tau$ is a (constant) relaxation time, and $v_{i}$ is the group velocity at energy $\epsilon_{i}$.
We then assembled feature vectors
$\mathbf{x}_{i}=\bigl{[}k_{ix},\,k_{iy},\,k_{iz},\,\epsilon_{i},\,n_{F}(\epsilon_{i})\bigr{]},$
and set $\sigma_{i}$ from Eq. (23) as the target variable.
Data Splitting and Normalization
The final dataset was split into training and testing subsets (e.g., 80% and 20%). We applied standard scaling (mean subtraction and variance normalization) to stabilize SVM training.
3.1.1 Comparison with the Kubo Formula
@@
Its emphasis lies in the derivative of $n_{F}$ near the Fermi level, reflecting that conduction arises from electrons whose occupancy changes significantly with small energy shifts.
By contrast, our SVM-based approach uses the occupancy directly, $\sigma_{i}=e^{2}v_{i}^{2}\,n_{F}(\epsilon_{i})\,\tau$, thus omitting the explicit derivative factor in Eq. (24). This simplification can still yield practical, accurate predictions for the system at hand, especially if $n_{F}(\epsilon)$ varies smoothly or if conduction is dominated by states near the Fermi energy.
3.1.2 Methodology
In the Code Availability section of this work, see svm_conductivity.py, we employ SVR to perform regression on band-gap and density features, aiming to predict a continuous conductivity value. Because SVR is the real-valued extension of SVM.
Kernel Selection.
We tested several kernel functions in $\mathrm{SVR}$ (Support Vector Regression):
  • RBF Kernel: Known for its Gaussian form, often mirroring Green’s function structures.
  • Polynomial Kernel: Introduces polynomial feature interactions.
  • Linear Kernel: Represents a simple linear hypothesis in feature space.
  • Sigmoid Kernel: Resembles neural net activation but may fail to ensure positive semi-definiteness under certain parameter settings.
Hyperparameter Tuning
We used cross-validation to optimize parameters such as $C$ (regularization), $\gamma$ (RBF or sigmoid width), and polynomial degree. The best configuration under each kernel was then retrained on the entire training set.
Performance Metrics
We used Mean Squared Error (MSE) and the coefficient of determination ($R^{2}$) to evaluate predictive accuracy on the test set.
3.1.3 Results and Observations
Figure 2 compares kernel performance in predicting $\sigma_{i}$ from real copper data. The RBF kernel outperforms polynomial, linear, and sigmoid kernels, highlighting its aptitude for capturing nonlinear, Gaussian-like relationships in electron conduction.
The impressive performance of the RBF kernel indicates that kernels approximating Gaussian or propagator shapes naturally suit electron conduction problems. Furthermore, even though our approach omits the derivative term in Kubo’s formula, the kernel delivered strong performance for copper, indicating it may be sufficient when $n_{F}(\epsilon)$ changes gradually near the Fermi level. Nonetheless, investigating partial-derivative-based features could further refine predictive accuracy.
Figure 2: Comparison of SVM kernel performance in predicting electrical conductivity. The RBF kernel performs well in modeling complex non-linearities, reflecting its alignment with the underlying physical (Green’s-function-like) processes. The sigmoid kernel performs poorly due to incompatibilities with these processes.
3.2 Scattering Amplitudes and Linear Kernel
3.2.1 Physical Background
@@
3.2.2 Simulating setup
Dataset Preparation
We simulated a dataset representing electron-proton scattering events. Each data point included:
  • Incident Energy $E_{i}$
  • Scattering Angle $\theta$
  • Scattering Amplitude $f(E_{i},\theta)$
Generating Scattering Data
@@ 3.2.2 Simulating setup
3.2.3 Low-Energy Domain and Code Implementation
The reference file scattering_linear_kernel.py for our scattering experiment can be found in Section Code Availability. In it, we consider frequencies in the range $\nu\approx(5\times 10^{14})\ldots(7\times 10^{14})\,\mathrm{Hz}$, focusing on the photoelectric effect near a work function of $\phi=2.0\,\mathrm{eV}$. In this low-energy regime, the scattering/emission process is adequately described by a simple linear relationship:
@@ 3.2.3 Low-Energy Domain and Code Implementation
where $E_{\mathrm{kin}}$ is the electron’s kinetic energy, $h$ is Planck’s constant, and $\nu$ is the photon frequency. These code-generated data reflect a near-threshold photoemission scenario rather than a high-energy or relativistic regime; consequently, the scattering amplitude effectively reduces to a linear function of $\nu$. This justifies our choice of the linear kernel in SVM: the underlying physics indicates a straightforward linear dependency on the input variable $\nu$, hence more complex kernels (e.g. RBF or polynomial) generally offer little improvement for this near-threshold photoemission range. Indeed, we do not attempt to simulate higher-energy collisions where strong nonlinear or relativistic effects become relevant. Therefore, the dataset in this paper comprises only such low-energy photon–electron interactions, aligned with the code’s simplified “scattering amplitude” approach.
3.2.4 Methodology
we demonstrate how to use an SVR model to regress the kinetic energy $E_{\mathrm{kin}}$ from the incident photon frequency $\nu$. We select a linear kernel to illustrate near-threshold behavior, evaluating MSE and $R^{2}$ on simulated data.
The feature vectors we choosen were assembled as $\mathbf{x}_{i}=[E_{i},\theta_{i}]$, and the target variable was $y_{i}=f(E_{i},\theta_{i})$.
@@ 3.2.4 Methodology
3.2.5 Results and Discussion
Green’s Function in Electron-Proton Scattering
@@ 3.2.5 Results and Discussion
where $\hat{H}$ is the Hamiltonian describing the electron–proton system. For sufficiently small momentum transfers and low incident energies, $\hat{H}$ can often be treated as an essentially linear operator in this regime; the resulting Green’s function thus behaves in a way that is well-approximated by plane-wave-like or linear expansions. In other words, the system’s scattering amplitude depends almost linearly on the input variables $(E_{i},\theta)$ in the first Born approximation.
Because of this linear behavior, an SVM linear kernel (instead of an RBF or other more complex kernel) can naturally match the underlying Green’s function form for low-energy scattering. Hence, selecting a linear kernel in SVM regression mirrors the dominantly linear Green’s function in such scattering processes.
Discussion
The linear kernel outperformed the others, aligning with the linear dependence of the scattering amplitude on the input variables in the first Born approximation. This suggests that for scattering processes with inherently linear relationships, simple linear kernels suffice, negating the need for more complex kernels like RBF or sigmoid.
3.3 Anharmonic Oscillator Energy Levels and Polynomial Kernel
3.3.1 Physical Background
@@
3.3.2 Methodology
In the Code Availability section of this work, see anharmonic_polynomial_kernel.py, we applies SVR to regress the continuous force $F(x)=kx+\alpha x^{3}$ from simulated displacement–force data. We compare multiple polynomial kernel degrees to find which best matches the nonlinearity of the oscillator. MSE and $R^{2}$ quantify performance.
@@ 3.3.2 Methodology
$\bigl{\{}(n_{i},\alpha_{i},\beta_{i}),\;E_{n_{i}}\bigr{\}}.$
Polynomial Kernels of Varying Degrees and the Anharmonic Green’s Function
Although the anharmonic potential includes an $x^{4}$ term, its exact Green’s function (i.e., the resolvent of $\hat{H}=-\frac{\hbar^{2}}{2m}\frac{d^{2}}{dx^{2}}+\frac{1}{2}kx^{2}+\frac{\alpha}{4}x^{4}$) is generally quite complicated to derive in closed form. Nonetheless, one can approximate this Green’s function via polynomial expansions or perturbation approaches. Such expansions often emphasize the dominant cubic nonlinearity (from the force $F(x)=kx+\alpha x^{3}$), especially if higher-order corrections are negligible in a certain parameter range.
@@
for degrees $d=2,3,4,\dots$ on the dataset $\{(\mathbf{x}_{i},\;E_{n_{i}})\}$, where $\mathbf{x}_{i}=[n_{i},\,\alpha_{i},\,\beta_{i}]$ encodes quantum number and parameters, and $E_{n_{i}}$ is the corresponding energy level. For each $d$, we used cross-validation to select optimal $(\gamma,r)$.
Our findings indicate that $d=3$ consistently yielded the lowest prediction error, in alignment with the system’s dominant cubic behavior. One might initially expect $d=4$ (matching the $x^{4}$ potential term) to excel, but in practice, the cubic polynomial kernel often captures the essential nonlinearity while avoiding overfitting. Degrees above 3 did not show improved accuracy and, in some trials, led to unstable training. This outcome suggests that even a partial approximation of the true Green’s function (via a degree-3 polynomial) can effectively model the key physics in this anharmonic-oscillator example.
In addition to testing polynomial kernels (with varying degrees), we also trained SVM regression models using linear, RBF, and sigmoid kernels for the anharmonic oscillator dataset. We then compared their mean squared error (MSE) and $R^{2}$ scores.
Comparison Among Different Polynomial Degrees
Figure 4: Comparison of polynomial kernels at different degrees ($2,3,4$) for the anharmonic oscillator. Degree 3 consistently yields the best predictive accuracy.
We also tested polynomial kernels with varying degree from $d=2$ through $d=4$. As shown in Figure 4, degree 3 consistently achieved the lowest MSE on the test set. Although one might expect $d=4$ to match the $x^{4}$ term in the potential, in practice $d=3$ effectively captured the essential anharmonicity while avoiding overfitting issues.
Comparison Among Different Kernel Types
Figure 5 compares four distinct kernel forms: RBF, linear, sigmoid, and a polynomial of degree 3. While the RBF kernel also captures nonlinearity, the degree-3 polynomial kernel gave the best average performance in this setting, presumably reflecting the cubic nonlinearity of the anharmonic force $F(x)=kx+\alpha x^{3}$.
Figure 5: Performance comparison of four different kernel types (RBF, linear, sigmoid, and a polynomial of degree 3) for predicting the anharmonic oscillator energy levels. The degree-3 polynomial kernel outperforms the others on average.
@@ 3.4 Photonic Crystals
3.4.1 Methodology
In the Code Availability section of this work, see custom_kernel.py, we implement a SVR class that leverages a user-defined periodic kernel to predict phonon dispersion $\omega(k)$, treating it as a continuous regression problem. The custom kernel encodes periodicity akin to a crystal structure.
We simulated the band structure of a one-dimensional photonic crystal. Each data point included:
  • Wave Vector $k$
  • Frequency $\omega(k)$
Calculating Band Structure
@@ 3.4.1 Methodology
In addition to the standard linear, polynomial, and RBF kernels, we also designed a custom kernel incorporating photonic-crystal periodicity (described below). Hence, we trained and compared four kernel variants:
  • Linear Kernel
  • Polynomial Kernel (e.g., degree 3 for example)
  • RBF Kernel
  • Custom Periodic Kernel
The results in Fig. 6 show that the RBF and the custom kernel exhibit superior performance (better MSE and $R^{2}$) compared to linear or polynomial in this photonic-crystal problem.
Custom Kernel Design
Photonic crystals have a periodic refractive index, which influences photon propagation similarly to how periodic potentials affect electrons. While deriving an explicit Green’s function for such complex structures is challenging, we can design custom kernels that approximate the expected behavior. Hence, in addition to the standard SVM kernels, we designed a custom kernel incorporating the periodicity of photonic crystals. The custom kernel is
@@
The kernel combines a Gaussian decay with a cosine term to incorporate both local similarity and periodicity. This design is inspired by the qualitative form a Green’s function for a photonic crystal might exhibit, capturing essential features such as periodicity and exponential localization without needing its explicit derivation.
3.4.2 Results and Discussion
@@
Although the custom periodic kernel was designed to capture the specific periodicity and symmetry of photonic crystals, our experimental results indicate that the RBF kernel performs comparably well, with slightly better predictive accuracy in this instance. This outcome suggests that while the custom kernel aligns closely with the physical structure of the system, the RBF kernel’s inherent Gaussian properties already capture much of the essential local behavior. The custom kernel, however, remains valuable as it incorporates explicit periodic features, which may prove advantageous in other scenarios or with further parameter tuning. Thus, the slight performance difference does not diminish the significance of designing kernels based on physical insights; instead, it highlights the RBF kernel’s robustness and encourages further refinement of custom kernels to surpass standard kernels consistently.
3.5 Discussion of results
While the mathematical equivalence proven in the theoretical sections states that, under ideal conditions, SVM kernels can equal Green’s functions, practical applications often use kernels that approximate the behavior of Green’s functions rather than computing them explicitly. In electron-proton scattering, the success of the RBF kernel reflects its similarity to the linear Green’s function of electron propagation. For the anharmonic oscillator, although the exact Green’s function is complex, the polynomial kernel captures the nonlinearity of the system in a way that aligns with the theoretical framework. In photonic crystals, the custom kernel mimics key features of the expected Green’s function - periodicity and decay - demonstrating how physical insight informs kernel design. These examples underscore that even if we do not compute explicit Green’s functions in each application, the selection and design of kernels are guided by the same principles, bridging theory and practice.
4 Beyond Standard Kernels: Addressing Sigmoid-Like Limitations through the Kernel Polynomial Method (KPM)
In this section, we address the challenges presented by non-positive semi-definite (non-PSD) kernels in Support Vector Machines (SVMs) by leveraging the Kernel Polynomial Method (KPM). Non-PSD kernels, such as sigmoid kernels, often fail to satisfy Mercer’s condition, which is critical for SVM optimization [13, 2]. This failure leads to non-convex optimization problems and unstable training. Existing approaches to handle non-PSD kernels include spectral corrections (e.g., clipping negative eigenvalues) [15, 16], difference of convex functions programming [17], and constructing Krein space embeddings [18].
Here, we propose an alternative approach by integrating KPM into SVM kernel design. Specifically, we utilize orthogonal polynomial expansions, such as Jacobi and Chebyshev polynomials, to approximate kernel functions and enforce positive semi-definiteness [26, 27]. This approach not only addresses non-PSD issues but also aligns kernel design with physical principles like Green’s functions, which naturally exhibit Gaussian-like decay [19, 20].
The proposed KPM-based framework has several key advantages. Firstly, KPM ensures that by directly modifying the spectral properties of the respective kernel matrix the constructed kernels satisfy Mercer’s condition. Secondly, KPM retains the original structure of the non-PSD kernels while ensuring positive semi-definiteness. And thirdly, for applications involving physical models, kernels based on KPM usually tend to better capture the underlying Green’s function properties compared to standard kernels.
4.1 Exceptions for SVM Kernels
The non-positive semidefinite (non-PSD) kernels include sigmoid kernels, indefinite similarity matrices (e.g., Smith-Waterman), and non-standard kernel functions. All of them can be used with SVMs, however the sigmoid kernel is by far the most common one.
Limitations of the Sigmoid Kernel

### v2

@@ 4 SVM regression in physical systems
In the following experiments, we predominantly utilize synthetic datasets. Real-world experimental data often contains inherent noise and measurement uncertainties. By using synthetic data generated from known physical equations (e.g., the first Born approximation, the anharmonic oscillator equation), we can isolate the effect of the kernel choice from other variables and therefore obtain a clearer confirmation for our hypothesis that kernel functions aligned with Green’s functions improve the predictive capability of a model.
4.1 Computing conductivity
Data acquisition.
For this study, we retrieve copper (Cu) data from the Materials Project database [29]. Following the procedures described in Section 4.1, we extract available band structures, Fermi levels, and $\mathbf{k}$-points relevant to copper-based materials entries. In principle, each $\mathbf{k}_{i}$ may correspond to an electronic energy $\epsilon_{i}$ in the band structure. However, for demonstration purposes in the present section, we omit partial-derivative features and instead construct a simpler dataset focusing only on the band gap $E_{g}$ and density $\rho$ of relevant copper entries.
Feature construction.
@@ 4.1 Computing conductivity
Its emphasis lies in the derivative of $n_{F}$ near the Fermi level, which is due to the fact that conduction arises from electrons whose occupancy changes significantly with small energy shifts.
Here however, we adopt a synthetic conductivity function to show how different SVM kernels behave in capturing non-linear relationships. Specifically, we define
$\sigma\;=\;\exp\bigl{(}-E_{g}\bigr{)}\,\sqrt{\rho},$ (28)
with the band gap $E_{g}$ and the material density $\rho$ as input features. This choice avoids potential numerical instabilities associated with partial derivatives of the Fermi function, and instead provides a smooth exponential–square-root relationship that more directly isolates kernel effects. While this is not a physical replacement for the Kubo approach, it serves to highlight how kernel choices can strongly influence regression outcomes, especially in capturing nonlinear dependencies.
We assembled a dataset of $(E_{g},\rho,\sigma)$ tuples for the experiment of $N=128$ points.
Data splitting and normalization.
Following standard practice, we partition our dataset into training and testing subsets (80% vs. 20%). Additionally, we apply zero-mean, unit-variance scaling to each feature dimension and to the target $\sigma$ to promote stable training. Such normalization mitigates the effects of disparate numerical ranges among $E_{g}$, $\rho$, and $\sigma$.
Training and hyperparameter tuning.
We trained support vector regression ($\mathrm{SVR}$) models under four different kernel types: RBF, polynomial with $\mathrm{degree}=3$, linear, and sigmoid. We used basic or grid-based searches to identify near-optimal $(C,\gamma)$ parameters for each kernel.
In principle, these hyperparameters can be further optimized via cross-validation to minimize mean squared error (MSE) on the training folds. For clarity, our comparisons simply fix a moderate $C$ (e.g. $C=100$) and $\gamma=\texttt{scale}$, since the primary goal is to observe the qualitative differences among kernels.
Results and analysis
Figure 2 presents a comparison of the mean squared error (MSE) and coefficient of determination ($R^{2}$) achieved by each kernel on the held-out test set. Panel (a) shows that the RBF kernel achieves the lowest MSE, which indicates its strength in capturing Gaussian-like nonlinearities. In contrast, the sigmoid kernel exhibits substantially larger MSE (and may even yield negative $R^{2}$), indicating poor compatibility under this parameterization. The polynomial kernel and the linear kernel each capture partial trends, but remain less accurate than RBF on the smoothly varying synthetic $\sigma$.
These findings echo the observations reported in more rigorous, physics-based modeling (see e.g. [31]), where partial-derivative-based features can reveal additional structures near the Fermi level, but also add numerical complexity. Hence, this simplified demonstration strongly suggests that an RBF kernel is a robust choice when dealing with conductivity-like regressions, especially if the underlying relationships are smooth exponentials or mild polynomials. Future work could incorporate physically grounded features (e.g. $v_{i}^{2}\,n_{F}(\epsilon_{i})$, or the Kubo derivative) to further refine predictive accuracy while retaining the kernel flexibility.
Figure 2: Comparison of four SVM kernels in predicting a synthetic conductivity. (a) Mean Squared Error (MSE) for each kernel. The RBF kernel achieves the lowest error, indicating strong ability to model nonlinear relationships; the sigmoid kernel yields a very high MSE, reflecting poor compatibility. We also compare a polynomial kernel with $\mathrm{degree}=3$, and a simpler linear kernel. (b) Coefficient of determination $R^{2}$ across the same four kernels, again highlighting the superior performance of the RBF kernel.
4.2 Scattering amplitudes
Physical background.
@@ 4.2 Scattering amplitudes
Dataset preparation.
We generated a synthetic dataset of 200 samples representing electron-proton scattering events. Each data point included incident energy $E_{i}$, scattering angle $\theta$, and scattering amplitude $f(E_{i},\theta)$.
@@
In our experiment, we consider frequencies in the range $\nu\approx(5\times 10^{14})\ldots(7\times 10^{14})\,\mathrm{Hz}$, focusing on the photoelectric effect near a work function of $\phi=2.0\,\mathrm{eV}$. In this low-energy regime, the scattering/emission process is adequately described by a simple linear relationship:
@@
where $E_{\mathrm{kin}}$ is the electron’s kinetic energy, $h$ is Planck’s constant, and $\nu$ is the photon frequency.
These synthetically generated data reflect a near-threshold photoemission scenario rather than a high-energy or relativistic regime. Consequently, the scattering amplitude effectively reduces to a linear function of $\nu$.
This indicates that the linear kernel would likely perform well for this problem. Consequently, more complex kernels (e.g. RBF or polynomial) are expected to generally offer little improvement for this near-threshold photoemission range.
Here, we do not attempt to simulate higher-energy collisions where strong nonlinear or relativistic effects become relevant. Therefore, the dataset in this paper comprises only such low-energy photon-electron interactions.
Training.
We demonstrate how to use an SVR model to regress the kinetic energy $E_{\mathrm{kin}}$ from the incident photon frequency $\nu$.
The feature vectors were assembled as $\mathbf{x}_{i}=[E_{i},\theta_{i}]$, and the target variable was $y_{i}=f(E_{i},\theta_{i})$.
@@
Results and analysis.
The results are shown in Fig. 3. The linear kernel outperformed the others, aligning with the linear dependence of the scattering amplitude on the input variables in the first Born approximation. This suggests that for scattering processes with inherently linear relationships, simple linear kernels suffice, negating the need for more complex kernels like RBF or sigmoid.
@@
where $\hat{H}$ is the Hamiltonian describing the electron–proton system. For sufficiently small momentum transfers and low incident energies, $\hat{H}$ can often be treated as an essentially linear operator in this regime. The resulting Green’s function therefore behaves in a way that is well-approximated by plane-wave-like or linear expansions. In other words, the system’s scattering amplitude depends almost linearly on the input variables $(E_{i},\theta)$ in the first Born approximation.
Because of this linear behavior, an SVM linear kernel can naturally match the underlying Green’s function form for low-energy scattering. mirrors the dominantly linear Green’s function in such scattering processes.
4.3 Energy levels of anharmonic oscillators
Physical background.
@@ 4.3 Energy levels of anharmonic oscillators
Dataset generation.
For this problem, we apply SVR to regress the continuous force in Eq. (32) from simulated displacement–force data. We compare multiple polynomial kernel degrees to find which best matches the nonlinearity of the oscillator. MSE and $R^{2}$ quantify performance.
@@
$\bigl{\{}(n_{i},\alpha_{i},\beta_{i}),\;E_{n_{i}}\bigr{\}}$
where $n_{i}$ is the quantum number. The number of samples in the dataset was 500.
Training.
Although the anharmonic potential includes an $x^{4}$ term, its exact Green’s function (i.e., the resolvent of $\hat{H}=-\frac{\hbar^{2}}{2m}\frac{d^{2}}{dx^{2}}+\frac{1}{2}kx^{2}+\frac{\alpha}{4}x^{4}$) is generally quite complicated to derive in closed form. Nonetheless, one can approximate this Green’s function via polynomial expansions or perturbation approaches. Such expansions often emphasize the dominant cubic nonlinearity from the force Eq. (32), especially if higher-order corrections are negligible in a certain parameter range.
@@
for degrees $d=2,3,4,\dots$ on the dataset $\{(\mathbf{x}_{i},\;E_{n_{i}})\}$, where $\mathbf{x}_{i}=[n_{i},\,\alpha_{i},\,\beta_{i}]$. For each $d$, we used cross-validation to select optimal $(\gamma,r)$.
In addition to testing polynomial kernels with varying degrees, we also trained SVM regression models using linear, RBF, and sigmoid kernels for the anharmonic oscillator dataset.
Results and analysis.
Figure 4: Comparison of polynomial kernels at different degrees (2,3,42,3,42 , 3 , 4) for the anharmonic oscillator. Degree 3 consistently yields the best predictive accuracy.
As shown in Figure 4, degree 3 polynomial kernel consistently achieved the lowest MSE on the test set, compared to other degrees. Although one might expect $d=4$ to match the $x^{4}$ term in the potential, in practice $d=3$ effectively captured the essential anharmonicity while avoiding overfitting issues. Degrees above 3 did not show improved accuracy and, in some trials, led to unstable training. This outcome suggests that even a partial approximation of the true Green’s function (via a degree-3 polynomial) can effectively model the key physics in this anharmonic-oscillator example.
Figure 5 compares four distinct kernel forms: RBF, linear, sigmoid, and a polynomial kernel of degree 3. While the RBF kernel also captures nonlinearity, the degree-3 polynomial kernel gave the best average performance in this setting, presumably reflecting the cubic nonlinearity of the anharmonic force $F$ in Eq. (32).
Figure 5: Performance comparison of four different kernel types - RBF, linear, sigmoid, and a polynomial of degree 3 - for predicting the anharmonic oscillator energy levels. The degree-3 polynomial kernel outperforms the others on average.
@@ 4.4 Photonic crystals
Dataset generation.
Here, we implement a SVR class that leverages a user-defined periodic kernel to predict phonon dispersion $\omega(k)$, treating it as a continuous regression problem. The custom kernel encodes periodicity akin to a crystal structure.
We simulate the band structure of a one-dimensional photonic crystal using 500 data points. Each data point included wave vector $k$, and frequency $\omega(k)$.
Calculating band structure.
@@
Custom kernel design.
Photonic crystals have a periodic refractive index, which influences photon propagation. Deriving an explicit Green’s function for such complex structures is challenging. Here, we instead design a custom kernel that approximates the expected behavior. Hence, in addition to the standard linear, polynomial, and RBF kernels, we also performed simulations with a custom kernel incorporating photonic-crystal periodicity.
The custom kernel is defined as
@@
The kernel combines a Gaussian decay with a cosine term to incorporate both local similarity and periodicity. This design is inspired by the form that a Green’s function for a photonic crystal might assume.
Results and analysis.
@@
Although the custom periodic kernel was designed to capture the specific periodicity and symmetry of photonic crystals, our experimental results indicate that the RBF kernel also performs comparably well, with slightly better predictive accuracy in this instance. This outcome suggests that while the custom kernel aligns closely with the physical structure of the system, the RBF kernel’s inherent Gaussian properties already capture much of the essential local behavior.
The custom kernel, however, remains valuable as it incorporates explicit periodic features, which may prove advantageous in other scenarios or with further parameter tuning. Thus, the slight performance difference does not diminish the significance of designing kernels based on physical insights. Instead, it highlights the RBF kernel’s robustness.
5 Custom kernel design with KPM
The non-PSD kernels include sigmoid kernels, indefinite similarity matrices (e.g., Smith-Waterman), and non-standard kernel functions. All of them can be used with SVMs, however the sigmoid kernel is by far the most common one.